[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Oct 31, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Oct 31 15:05:51 UTC 2017


Here is today's summary of NHC computer operations:
--- Craig Mattocksi

----------

TAFB:
1. 00Z NWPS data was never received for NH2 despite CheckNWPS message  
showing it was running and the WRKNWA message showing it was complete. NCO  
Sysadmin Doug Gaer checked - the NWPS run finished normally and the output  
looks fine. TSB AWIPS II Sysadmin Chris Mello will check whether there was  
a problem with the EDEX data ingest last night.

2. Collaboration server problems (reprise) - Chat server still occasionally  
kicks users off the LX5 AWIPS II workstation. Chris Mello will check  
whether NCF Trouble Ticket #152191 has been elevated to HIGH priority.

Update:
Doug Gaer recently rebooted the comp02 server, which has solved many  
problems with the NWPS. TAFB Forecasters, please monitor.
HSU:
Update:
Matt Sardi discovered that NHC has never ingested or decoded data from  
Pulaski Shoals. He has added this C-MAN station to our list of stations to  
be docoded for display in N-AWIPS.
Other items:
Update:
William Martinez is investigating problems with the AC and temperature  
fluctuations in the server room.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171031/3cccc679/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list